# A Survey on the Safety and Security Threats of Computer-Using Agents: JARVIS or Ultron?

### v2

@@ 5 Evaluation and Benchmarking
Computer-Using agents (CUAs) are widely deployed across various platforms, necessitating a more comprehensive evaluation of their safety performance compared to general LLM-based agents. This section provides a structured summary of these benchmarks, as shown in Table 4.
@@ 5.1.1 Web-based Scenario
In the web-based scenario, several datasets have been proposed to assess the safety of agents operating within browser environments. Specifically, ST-WebAgentBench (Levy et al., 2024) and BrowserART (Kumar et al., 2024) focus on evaluating agents’ safety-related behaviors in tasks involving web navigation, interaction, and tool usage under potential prompt injection threats. Meanwhile, PrivacyLens (Shao et al., 2024) investigates privacy-sensitive interactions in web-based conversations, containing 493 validated prompts derived from U.S. legal, social, and interpersonal communication norms. In parallel, CASA (Qiu et al., 2025) provides a web-based benchmark designed to evaluate agents’ awareness of cultural and social contexts, utilizing grounded questions and descriptors sourced from CultureBank. Furthermore, ShieldAgent-Bench (Chen et al., 2025c) extends these efforts by simulating adversarial instructions and policy-violation scenarios across diverse web environments, providing 960 safety-related instructions and 3,110 unsafe trajectories. Finally, the VWA-Adv benchmark (Wu et al., 2024a) targets web-based scenarios, introducing 200 adversarial tasks built on VisualWebArena (Koh et al., 2024) to evaluate agent robustness against realistic attacks through imperceptible webpage perturbations and component-wise adversarial flows.
@@ 5.1.2 Mobile-based Scenario
MobileSafetyBench (Lee et al., 2024a) targets the mobile environment, where agents are tested with real mobile applications. The dataset consists of 80 representative tasks across domains such as messaging, social media, finance, and system utilities. It serves to evaluate agents’ safety performance under mobile-specific constraints and risks.
@@ 5.1.3 General-purpose Scenario
Prompt injection remains a major concern for agent safety. InjecAgent (Zhan et al., 2024) and AgentDojo (Debenedetti et al., 2024) target this threat, testing agents equipped with diverse toolsets and under various attacker scenarios. InjecAgent features 330 tools from 36 distinct toolkits, while AgentDojo includes 97 realistic tasks and 629 security-focused test cases. The AgentHarm dataset (Andriushchenko et al., 2024) generalizes harmful behavior testing, introducing 110 base behaviors grouped into 11 harm categories such as fraud, cybercrime, and misinformation.
Besides, PrivacyLens (Shao et al., 2024) generalizes privacy risk evaluation, introducing 493 privacy-sensitive vignettes and trajectories to assess information leakage in language models across various scenarios. In addition, Agent-SafetyBench (Zhang et al., 2024e) builds upon these efforts by refining general-purpose samples to improve diversity and cover underrepresented risk types. Furthermore, the Agent Security Bench (ASB) (Zhang et al., 2024b) serves as a comprehensive and large-scale benchmark, encompassing 10 distinct scenarios, 10 purpose-built agents, and over 400 tools and tasks. ASB is designed to provide a unified framework for evaluating the security of LLM-based agents across a wide range of operational contexts.
@@ 5.2.1 Real-world Environments
Several studies have focused on evaluating agent behavior in such contexts. ST-WebAgentBench (Levy et al., 2024), BrowserART (Kumar et al., 2024), and CASA (Qiu et al., 2025) evaluate agent behavior on real websites, assessing aspects such as safety, trustworthiness, and cultural awareness. Unlike prior benchmarks that only provide task instructions, ShieldAgent-Bench (Chen et al., 2025c) evaluates real-world safety challenges by incorporating complete agent interaction protocols, such as instructions, trajectories, enforced policies, and ground-truth labels. Similarly, MobileSafetyBench (Lee et al., 2024a) examines agent performance across various mobile applications, including messaging, web navigation, social media, finance, and utility apps. To provide a more comprehensive assessment, Agent-SafetyBench (Zhang et al., 2024e) extends evaluations across both web and mobile environments, offering a broader analysis of agent safety. In contrast, the VWA-Adv benchmark (Wu et al., 2024a) focuses on a realistic threat model in real-world web environments, where the attacker is a legitimate user with limited capabilities to manipulate the environment.
In the context of security vulnerabilities introduced by external content, InjecAgent (Zhan et al., 2024) simulates agent responses to adversarial inputs, enabling researchers to assess the agent’s subsequent actions and obtain attack outcomes. PrivacyLens (Shao et al., 2024) analyzes the data transmitted by agents to external tools, identifying potential privacy risks during task execution.
@@ 5.2.2 Sandbox Environments
R-Judge (Yuan et al., 2024), TrustAgent (Hua et al., 2024), AgentDojo (Debenedetti et al., 2024), and AgentHarm (Andriushchenko et al., 2024) each construct a simulation environment with distinct focuses. R-Judge adopts ReAct (Yao et al., 2023) as its interactive framework. TrustAgent enhances task evaluation by providing detailed descriptions of external tools relevant to each task domain. AgentDojo builds an extensible environment designed for developing and assessing new agent tasks, defenses, and adaptive attacks. Meanwhile, AgentHarm utilizes synthetic tools for all tasks, implemented through Inspect (AI Security Institute, ), effectively mimicking a range of tools from general utilities to domain-specific applications.
@@ 5.3.1 Task Completion Metrics
Task Success Rate (TSR) assesses whether an agent successfully reaches the final goal of a task, regardless of the performance on intermediate steps (Yao et al., 2022; Xie et al., 2024; Wen et al., 2023). It serves as a holistic indicator of an agent’s overall effectiveness in completing a given task. In many safety benchmarks and datasets, this measure is akin to the Benign Success Rate (Benign SR) (Wu et al., 2024a) or Benign Utility (Debenedetti et al., 2024) or Performance Under No Attack (PNA) (Zhang et al., 2024b), which evaluates how well an agent performs under normal, non-adversarial conditions. A high Task Success Rate, therefore, not only signifies that the agent meets the intended outcome but also underlines its reliability in standard operational settings.
@@ 5.3.3 Safety and Robustness Metrics
Similarly, the Violation Rate (Vio-R) in CASA (Qiu et al., 2025) measures the fraction of agent responses that violate stated norms when presented with misleading or malicious inputs, which effectively captures how often the agent is “attacked” into norm-breaking behavior.
@@
In addition to ASR, some benchmarks go beyond a binary success measure to characterize how severe or how partial an unsafe outcome is: In TrustAgent (Hua et al., 2024), the Safety metric evaluates both the likelihood and the severity of potential risks by assigning categorical ratings (e.g., “Certain No Risk” to “Likely Severe Risk”) to each agent response, thus capturing not only whether a response is unsafe but how risky it is. In contrast, Agent-SafetyBench (Zhang et al., 2024e) simplifies the assessment with a Safety Score, measured as the proportion of test cases labeled “safe” by an LLM judge. Meanwhile, in AgentHarm (Andriushchenko et al., 2024), the Harm Score is computed via a detailed manually written grading rubric where outputs earn partial credit whenever some but not all harmful criteria are triggered, providing a finer-grained view of measure of how much of the harmful behavior was actually executed.
@@
a higher RR indicates that the agent is more cautious and conservative, often rejecting potentially harmful or policy-violating requests. However, if the agent exhibits a high RR even on low-risk or benign tasks, it may be unnecessarily refusing requests. In contrast, a lower RR means the agent is more permissive, which can enhance user experience but might also lead to a greater risk of unsafe outcomes (Lee et al., 2024a).
@@
This approach typically involves a classifier or human annotators to determine whether leaked information appears in each output.
@@
7 Inference Cost refers to the computational and monetary resources required to generate model outputs during deployment. In ShieldAgentBench (Chen et al., 2025c), it is quantified by the average number of closed-source LLM API queries and the end-to-end response time (in seconds) per sample, measured from the system runtime logs. Including inference cost alongside accuracy and robustness metrics offers a more complete view of an agent’s real-world performance and efficiency trade-offs.
@@ 5.4.1 Rule-based Measurements
Rule-based measurement involves the use of predefined rules or algorithms to compute evaluation metrics without manual annotation or LLM intervention. Typically implemented in code, these rules automatically assess agent behavior against fixed, deterministic criteria, making this approach suitable for well-defined and objective evaluation standards.
This method is widely adopted across existing agent safety benchmarks. For instance, ShieldAgent (Chen et al., 2025c) adopts this approach to directly compute evaluation metrics. TrustAgent (Hua et al., 2024) measures the overlap of action trajectories to assess goal alignment and safety compliance. AgentDojo (Debenedetti et al., 2024) and InjecAgent (Zhan et al., 2024) compute ASR variants using predefined criteria to capture attack effectiveness and resilience. PrivacyLens (Shao et al., 2024) uses binary (yes/no) rule-based judgments on privacy-sensitive prompts. Likewise, ST-WebAgentBench (Levy et al., 2024) applies programmatic functions to evaluate policy compliance via DOM and action traces. Both MobileSafetyBench (Lee et al., 2024a) and Agent-SafetyBench (Zhang et al., 2024e) rely on rule-based checks for task success and harm prevention, while Agent Security Bench (ASB) (Zhang et al., 2024b) adopts rule-based ASR calculations to quantify attack impact. In addition, AgentHarm (Andriushchenko et al., 2024) also employs predefined rules to evaluate most simple tasks, thereby minimizing dependence on LLM-based grading. Ultimately, VWA-Adv benchmark (Wu et al., 2024a) uses a rule-based approach to evaluate agent robustness, which models the agent as a directed graph and calculating adversarial influence along edges.
While rule-based methods are efficient, automated, and reproducible—enabling scalable evaluation—they lack the flexibility to handle nuanced or context-dependent agent behaviors.

### v3

@@ 5 Evaluation and Benchmarking
Computer-Using agents (CUAs) are widely deployed across various platforms, necessitating a more comprehensive evaluation of their safety performance compared to general LLM-based agents. This section provides a structured summary of these benchmarks, as shown in Table 4 and  5.
@@ 5.1.1 Web-based Scenario
In the web-based scenario, several datasets have been proposed to assess the safety of agents operating within browser environments. Specifically, ST-WebAgentBench (Levy et al., 2024) and BrowserART (Kumar et al., 2024) focus on evaluating agents’ safety-related behaviors in tasks involving web navigation, interaction, and tool usage under potential prompt injection threats. Meanwhile, PrivacyLens (Shao et al., 2024) investigates privacy-sensitive interactions in web-based conversations, containing 493 validated prompts derived from U.S. legal, social, and interpersonal communication norms. In parallel, CASA (Qiu et al., 2025) provides a web-based benchmark designed to evaluate agents’ awareness of cultural and social contexts, utilizing grounded questions and descriptors sourced from CultureBank. Furthermore, ShieldAgent-Bench (Chen et al., 2025e) extends these efforts by simulating adversarial instructions and policy-violation scenarios across diverse web environments, providing 960 safety-related instructions and 3,110 unsafe trajectories. SafeArena (Tur et al., 2025) likewise broadens coverage by injecting jailbreak-inspired malicious intents into 4 realistic WebArena sites and introducing 500 paired safe vs. harmful tasks over 5 harm categories. Similarly, WASP (Evtimov et al., 2025) combines 21 concrete attacker goals with 2 benign user goals under both URL and plaintext injection templates, with total of 84 tasks, to evaluate agent security against prompt injection attacks. VPI-Bench (Cao et al., 2025) targets visual prompt injection, providing 306 test cases across five popular sites, each embedding an adversarial instruction directly in the on-screen UI to see whether agents follow it. AgentDAM (Zharmagambetov et al., 2025) assesses AI agents’ propensity to expose sensitive information across three realistic web settings (Reddit, GitLab, Shopping) over 246 tasks. Finally, the VWA-Adv benchmark (Wu et al., 2024a) targets web-based scenarios, introducing 200 adversarial tasks built on VisualWebArena (Koh et al., 2024) to evaluate agent robustness against realistic attacks through imperceptible webpage perturbations and component-wise adversarial flows.
@@ 5.1.2 Mobile-based Scenario
Mobile-focused benchmarks evaluate CUAs inside real mobile applications. For instance, MobileSafetyBench (Lee et al., 2024a) consists of 80 representative tasks across domains such as messaging, social media, finance, and system utilities. It serves to evaluate agents’ safety performance under mobile-specific constraints and risks. Hijacking JARVIS (Liu et al., 2025a) contributes a two-part mobile benchmark: 58 dynamic tasks with attack patterns in different complexity levels and a 210-screenshot static set across 14 popular Android apps, enabling both live and offline safety evaluation.
@@ 5.1.3 General-purpose Scenario
Several datasets are designed with general-purpose safety evaluation in mind, spanning diverse tools, risks, and interaction environments.
Tool-use scenario. Tool-enabled CUAs have received intense scrutiny over the past two years. ToolEmu (Ruan et al., 2023) probes safety failures in a fully LM-emulated sandbox, covering 36 toolkits (18 categories), 144 high-stakes tasks, and 9 risk types. RAS-Eval (Fu et al., 2025) standardizes security testing for tool-driven agents with 80 core test cases and 3,802 attack tasks mapped to 11 CWE categories across both simulated and real tool executions. Prompt-injection–oriented suites such as AgentDojo (Debenedetti et al., 2024) with 97 realistic tasks, 629 security cases and InjecAgent (Zhan et al., 2024) with 330 tools from 36 toolkits evaluate how well agents perform under various adversarial scenarios while equipping with diverse tools. AgentHarm (Andriushchenko et al., 2024) broadens harmful-behavior testing with 110 base behaviors in 11 harm categories, and the large-scale Agent Security Bench (ASB) (Zhang et al., 2024b) aggregates 10 scenarios, 10 purpose-built agents, and over 400 tools and tasks to offer a unified safety framework. Furthermore, Agent-SafetyBench (Zhang et al., 2024e) covers 349 interaction environments and 2,000 test cases, spanning 8 safety risk categories and 10 prevalent failure modes in unsafe agent behaviors.
Mixed / hybrid environments Several benchmarks test agents that operate across heterogeneous interfaces (web, OS, shells, code executors, etc.). For instance, OpenAgentSafety (Vijayvargiya et al., 2025) provides 350 multi-turn, multi-user tasks in both benign and adversarial settings using a real browser, shell, file system, and messaging APIs. RiOSWorld (Yang et al., 2025b) runs 492 risky tasks in 13 categories on an OSWorld VM, capturing both environment and user-originated risks. RedTeamCUA (Liao et al., 2025) introduces RTC-Bench with 864 hybrid Web–OS adversarial scenarios, underscoring CUAs’ susceptibility to indirect prompt injection. MLA-Trust (Yang et al., 2025c) evaluates 34 high-risk, real-world tasks, showing how multi-step interactions in real environments can amplify risks beyond static LLM outputs. GUI-Robust (Yang et al., 2025a) complements this by injecting seven classes of interface anomalies (e.g., ad pop-ups, loading delays) to study robustness. Finally, HAICOSYSTEM (Zhou et al., 2024) emulates realistic human–AI interactions and complex tool use by running 8K+ simulations across 132 scenarios in seven domains, covering multi-dimensional risks (operational, content, societal, legal).
Broader risk-awareness and multidimensional safety. Beyond concrete tool or environment settings, several works emphasize comprehensive risk taxonomies and analysis. R-Judge (Yuan et al., 2024) scores risk awareness over 569 multi-turn interactions spanning 5 categories, 27 scenarios, and 10 risk types. TrustAgent (Hua et al., 2024) contributes 70 samples across 5 domains with paired ground-truth implementations to evaluate both helpfulness and safety. PrivacyLens (Shao et al., 2024) offers 493 privacy-sensitive vignettes and trajectories for leakage analysis. GUI-Robust (Yang et al., 2025a) complements these efforts by focusing on robustness under anomalies in interactions. It includes seven categories of common interface failures, such as advertisement pop-up and page loading delay.
@@ 5.2.1 Real-world Environments
Several studies have focused on evaluating agent behavior in such contexts. ST-WebAgentBench (Levy et al., 2024), BrowserART (Kumar et al., 2024), and CASA (Qiu et al., 2025) evaluate agent behavior on real websites, assessing aspects such as safety, trustworthiness, and cultural awareness. Unlike prior benchmarks that only provide task instructions, ShieldAgent-Bench (Chen et al., 2025e) evaluates real-world safety challenges by incorporating complete agent interaction protocols, such as instructions, trajectories, enforced policies, and ground-truth labels. Similarly, MobileSafetyBench (Lee et al., 2024a) examines agent performance across various mobile applications, including messaging, web navigation, social media, finance, and utility apps. Hijacking JARVIS (Liu et al., 2025a) also runs on real Andriod apps inside an emulator to create realistic on-screen manipulations, testing how agents behave under live, adversarial mobile interfaces. To provide a more comprehensive assessment, Agent-SafetyBench (Zhang et al., 2024e) extends evaluations across both web and mobile environments, offering a broader analysis of agent safety. OpenAgentSafety (Vijayvargiya et al., 2025) also evaluates agents in mixed live tool environments, automating a real browser, shell, file system, code executor, and messaging API. RAS-Eval (Fu et al., 2025) likewise executes attacks against live tools (e.g., MCP/JSON services) to measure real-world security failures, complementing its simulated setting. AgentDAM (Zharmagambetov et al., 2025) also operates in truly web-native settings, building atop the WebArena (Zhou et al., 2023) and VisualWebArena (Koh et al., 2024) simulators.
In contrast, the VWA-Adv benchmark (Wu et al., 2024a) focuses on a realistic threat model in real-world web environments, where the attacker is a legitimate user with limited capabilities to manipulate the environment. Likewise, VPI-Bench builds fully interactive, pseudo-authentic webpage replicas of selected websites—cloned DOMs rendered in a real browser—into which the authors manually inject the visual prompts, yielding a realistic yet controllable setting for testing on-screen attacks. Similarly, MLA-Trust (Yang et al., 2025c) evaluates CUAs in live GUI environments, including both websites and mobile apps, emphasizing how autonomous execution and dynamic interaction cycles can introduce compounding trust failures.
In the context of security vulnerabilities introduced by external content, InjecAgent (Zhan et al., 2024) simulates agent responses to adversarial inputs, enabling researchers to assess the agent’s subsequent actions and obtain attack outcomes. GUI-Robust (Yang et al., 2025a) similarly adopts real application interfaces with realistic anomalies, collecting human interaction traces to evaluate agent robustness under more complex or compound failure modes. PrivacyLens (Shao et al., 2024) analyzes the data transmitted by agents to external tools, identifying potential privacy risks during task execution.
@@ 5.2.2 Sandbox Environments
R-Judge (Yuan et al., 2024), TrustAgent (Hua et al., 2024), AgentDojo (Debenedetti et al., 2024), and AgentHarm (Andriushchenko et al., 2024) each construct a simulation environment with distinct focuses. R-Judge adopts ReAct (Yao et al., 2023) as its interactive framework. TrustAgent enhances task evaluation by providing detailed descriptions of external tools relevant to each task domain. AgentDojo builds an extensible environment designed for developing and assessing new agent tasks, defenses, and adaptive attacks. ToolEmu (Ruan et al., 2023) adopts a fully LM-emulated tool environment: instead of calling real APIs, a language model “pretends” to be each tool by returning text descriptions of what would have happened and updating a textual state. SafeArena (Tur et al., 2025) leverages four realistic WebArena sandbox sites (forum, e-commerce, code repo, retail), each instrumented per WebArena templates. RAS-Eval (Fu et al., 2025) also offers a controlled, tool-API sandbox, which simulates tool execution by letting tools read/write a shared in-memory state (e.g., calendars, databases). Likewise, by restricting attacker capabilities to creating issues, posts, and comments, WASP (Evtimov et al., 2025) offers a controlled yet realistic platform for end-to-end prompt injection testing.
Meanwhile, AgentHarm utilizes synthetic tools for all tasks, implemented through Inspect (AI Security Institute, ), effectively mimicking a range of tools from general utilities to domain-specific applications. RedTeamCUA (Liao et al., 2025) also constructs a hybrid sandbox by combining OS with Docker-based web platforms, enabling fully controlled realistic evaluation of indirect prompt injection across both web and OS interfaces. RiOSWorld (Yang et al., 2025b) runs agents inside an OSWorld virtual machine, exercising real desktop and web apps but under a fully instrumented, reproducible sandbox. HAICOSYSTEM (Zhou et al., 2024) also builds a fully controlled ecosystem—LLM-simulated users, tools, and environments inside a sandboxed OS/Web setting—enabling reproducible risk injections without live-site volatility.
@@ 5.3 Evaluation Metrics
This subsection summarizes the commonly used evaluation metrics for CUA safety. Most metrics reported in Table 4 and  5 are defined here. Extremely specialized metrics that are unique to a single work are omitted for brevity.
@@ 5.3.1 Task Completion Metrics
Task Success Rate (TSR) assesses whether an agent successfully reaches the final goal of a task, regardless of the performance on intermediate steps (Yao et al., 2022; Xie et al., 2024; Wen et al., 2023). It serves as a holistic indicator of an agent’s overall effectiveness in completing a given task. In many safety benchmarks and datasets, this measure is akin to the Benign Success Rate (Benign SR) (Wu et al., 2024a) or Benign Utility (Debenedetti et al., 2024) or Performance Under No Attack (PNA) (Zhang et al., 2024b), which evaluates how well an agent performs under normal, non-adversarial conditions.
Similarly, many benchmarks also measure how well agents complete required tasks, often including both benign and adversarial tasks. AgentDAM (Zharmagambetov et al., 2025) uses Utility measures how well agent performs its presrcibed tasks. HAICOSYSTEM (Zhou et al., 2024) uses Goal Completion (GOAL), scoring whether the agent achieves the scenario objective. In SafeArena (Tur et al., 2025), the Task Completion Rate (TCR) automatically matching each agent run against predefined reference objects to determine successful task completion. While, in RAS-Eval (Fu et al., 2025), TCR is the fraction of tasks where the agent’s tool-call sequence fully contains the human-annotated reference sequence; Task Incompletion Rate (TIR) counts runs that invoke only a subset or wrong combination of required tools; Task Fail Rate (TFR) flags runs that crash, call no tools, or exceed stack/length limits; and ultimately synthsizes these three metrics into a unified Performance Score (score).
@@ 5.3.3 Safety and Robustness Metrics
Since attacks vary in form and impact, several works define complementary metrics beyond ASR to capture different facets of adversarial effect. For instance, the Violation Rate (Vio-R) in CASA (Qiu et al., 2025) measures the fraction of agent responses that violate stated norms when presented with misleading or malicious inputs, which effectively captures how often the agent is “attacked” into norm-breaking behavior. Similarly, Misguided Rate (MisRate) in MLA-Trust (Yang et al., 2025c) measures how often an agent gives plausible yet incorrect answers to ambiguous or underspecified prompts. Formally, $\text{MisRate}=\frac{n_{\text{misguide}}}{n_{\text{amb}}}$, where $n_{\text{misguide}}$ is the number of misguided responses and $n_{\text{amb}}$ is the total count of ambiguous prompts. Likewise, Misleading Rate (MR) (Liu et al., 2025a) measures the fraction of attack trials in which the agent abandons the intended behavior and instead follows the injected misleading content. Risk Goal Completion (RGC) in RiOSWorld (Yang et al., 2025b) mirrors ASR by reporting the fraction of risky tasks where the agent fully achieves the attacker’s objective. In OS-Harm (Kuntz et al., 2025), a straightforward percentage of Unsafe execution traces has been calculated to quantify the overall safety-violation rate.
@@
In addition to ASR, some benchmarks go beyond a binary success measure to characterize how severe or how partial an unsafe outcome is: In TrustAgent (Hua et al., 2024), the Safety metric evaluates both the likelihood and the severity of potential risks by assigning categorical ratings (e.g., “Certain No Risk” to “Likely Severe Risk”) to each agent response, thus capturing not only whether a response is unsafe but how risky it is. ToolEmu (Ruan et al., 2023) also uses Safety with the similar idea: its safety evaluator flags failures and gives a 0–3 score by crossing harm likelihood with harm severity. In contrast, Agent-SafetyBench (Zhang et al., 2024e) simplifies the assessment with a Safety Score, measured as the proportion of test cases labeled “safe” by an LLM judge. Meanwhile, in AgentHarm (Andriushchenko et al., 2024), the Harm Score is computed via a detailed manually written grading rubric where outputs earn partial credit whenever some but not all harmful criteria are triggered, providing a finer-grained view of measure of how much of the harmful behavior was actually executed. HAICOSYSTEM (Zhou et al., 2024) refines the overall safety severity into five risk dimensions: Targeted Safety Risk (TARG), System/Operational Risk (SYST), Content Safety Risk (CONT), Societal Risk (SOC), Legal/ Rights Risk (LEGAL), each scored on a 10-10 to 0 scale to indicate how severely a trajectory violates that category.
@@
a higher RR indicates that the agent is more cautious and conservative, often rejecting potentially harmful or policy-violating requests. However, if the agent exhibits a high RR even on low-risk or benign tasks, it may be unnecessarily refusing requests. In contrast, a lower RR means the agent is more permissive, which can enhance user experience but might also lead to a greater risk of unsafe outcomes (Lee et al., 2024a). MLA-Trust (Yang et al., 2025c) instantiate RR as a Refusal-to-Execute Rate (RtE), where each output is judged (refuse / not refuse) by a specialized model such as GPT-4 or Longformer, following validated labeling protocols.
@@
AgentDAM (Zharmagambetov et al., 2025) likewise employs the LR metric to quantify any instances where sensitive data appears in an agent’s action outputs.
@@
7 Inference Cost refers to the computational and monetary resources required to generate model outputs during deployment. In ShieldAgentBench (Chen et al., 2025e), it is quantified by the average number of closed-source LLM API queries and the end-to-end response time (in seconds) per sample, measured from the system runtime logs. Including inference cost alongside accuracy and robustness metrics offers a more complete view of an agent’s real-world performance and efficiency trade-offs.
@@
9 Attempt Rate (AR) captures how often an agent attempts to follow an adversarial instruction, even if it never finishes the harmful task. In both RedTeamCUA (Liao et al., 2025) and VPI-Bench (Cao et al., 2025) LLM judges scrutinize the trajectory to flag any onset of unsafe behavior—RedTeamCUA uses a single LLM to detect beginnings of harmful actions, while VPI-Bench employs a majority vote of three frontier LLMs to decide whether an attack was “attempted”. It is also termed Risk Goal Intention (RGI) with the same idea in (Yang et al., 2025b).
10 Toxicity Score (TS) estimates how likely a response contains toxic, offensive, or harmful content by assigning a scalar value. (Yang et al., 2025c)
@@ 5.4.1 Rule-based Measurements
Rule-based measurement involves the use of predefined rules or algorithms to compute evaluation metrics without manual annotation or LLM intervention (Luo et al., 2025a). Typically implemented in code, these rules automatically assess agent behavior against fixed, deterministic criteria, making this approach suitable for well-defined and objective evaluation standards.
This method is widely adopted across existing agent safety benchmarks. For instance, ShieldAgent (Chen et al., 2025e) adopts this approach to directly compute evaluation metrics. TrustAgent (Hua et al., 2024) measures the overlap of action trajectories to assess goal alignment and safety compliance. AgentDojo (Debenedetti et al., 2024) and InjecAgent (Zhan et al., 2024) compute ASR variants using predefined criteria to capture attack effectiveness and resilience. PrivacyLens (Shao et al., 2024) uses binary (yes/no) rule-based judgments on privacy-sensitive prompts. Likewise, ST-WebAgentBench (Levy et al., 2024) applies programmatic functions to evaluate policy compliance via DOM and action traces. Both MobileSafetyBench (Lee et al., 2024a), Agent-SafetyBench (Zhang et al., 2024e) and WASP (Evtimov et al., 2025) rely on rule-based checks for task success, and even for harm prevention, while Agent Security Bench (ASB) (Zhang et al., 2024b) adopts rule-based ASR calculations to quantify attack impact. In addition, AgentHarm (Andriushchenko et al., 2024) also employs predefined rules to evaluate most simple tasks, thereby minimizing dependence on LLM-based grading. Furthermore, SafeArena (Tur et al., 2025) implements rule-based task-completion checks by matching agent outputs against predefined reference objects and uses the Agent Risk Assessment (ARIA) framework’s four hierarchical risk rules to quantify harmful-task outcomes. OpenAgentSafety (Vijayvargiya et al., 2025) implements Python-based evaluators that inspect the final environment state to detect unsafe outcomes. MLA-Trust (Yang et al., 2025c) adopts keywords matching method to automatically compute Refusal Rate. RAS-Eval (Fu et al., 2025) compares each agent’s tool-call sequence against a human-annotated reference sequence to automatically compute Task Completion, Incompletion, and Fail rates. RiOSWorld (Yang et al., 2025b) implement per-risk, rule-based evaluators that compare the final executable state to the expected unsafe outcome. Hijacking JARVIS (Liu et al., 2025a) evaluates the metrics with deterministic checks against task goals and labeled target elements with human annotated ground truth for the static set. AgentDAM (Zharmagambetov et al., 2025) also automatically inspects the final environment state against predefined success criteria to deterministically judge whether the agent completed its task. GUI-Robust (Yang et al., 2025a) evaluates agents by ensuring alignment with ground-truth trajectories throughout the task. Ultimately, VWA-Adv benchmark (Wu et al., 2024a) uses a rule-based approach to evaluate agent robustness, which models the agent as a directed graph and calculating adversarial influence along edges.
Rule-based approach can also be seen in some safety and security evaluation frameworks. For instance, the DoomArena framework (Boisvert et al., 2025) provides a modular library of scripted attack scenarios and built-in, deterministic checks on the final environment state, enabling automated detection of unsafe outcomes across a wide variety of web, OS, and multi-agent challenges.
While rule-based methods are efficient, automated, and reproducible—enabling scalable evaluation—they lack the flexibility to handle nuanced or context-dependent agent behaviors, making it hard to account for diverse unsafe attempts.